# New simple and accurate approximations for quintessence

### v2

@@ A.2 Tracking quintessence
Here we consider the tracking quintessence approximations given by Chiba (2010) [54], which in turn relied on earlier work by Watson and Scherrer (2003) [60], for the $\Gamma=\mathrm{constant}$ case, and, in particular, $\Gamma=\mathrm{constant}>1$, which yields the inverse power law potential. This work is based on a linearized equation for a perturbation of $w_{\phi}$ with respect to its asymptotic value $w_{\infty}$. To obtain an improved approximation for the linearized solution for $w_{\phi}$ Chiba [54] considered a series expansion in $e^{-3wN}$, when expressed in $e$-fold time and the present notation. This is the same series expansion quantity as in the present paper, but in contrast to Chiba [54] we used it to produce approximations for the full non-linear equations; thus, the expansions in the present work coincide with those in [60] and [54] for the quantities that have been computed in those two references up to first order but not for higher orders since Chiba [54] only deals with the linearized equation for $w_{\phi}$. Moreover, there is translation freedom in $N$. In the present work this freedom was fixed by the $[1/1]$ Padé approximant for $\Omega_{\phi}$, which resulted in an expansion in $T=\Omega_{\phi 0}e^{-3w_{\infty}N}/(1-\gamma\Omega_{\phi 0})$; Chiba [54] on the other hand fixed this freedom by, in effect, setting $w_{\phi}=w_{\infty}$, to obtain the following approximation for $\Omega_{\phi}$:
@@
which in turn yields an approximation for $E_{\phi}$, while the series expansion in $\tilde{T}$ for the linearized equation for $w_{\phi}$ resulted in
@@
Note that this is an expansion that is designed to improve the solution for the linear equation for $w_{\phi}$, not the actual solution of the non-linear equation for $w_{\phi}$. The improved linear approximation only improves the actual solution to the extent the linear approximation describes that solution, which, fortunately, to some extent is the case for the inverse power law potential, see e.g., Figure 1 in [28]. Not surprisingly, these approximations yield poorer accuracy than the present ones that are based on approximating the full non-linear equations.
@@ Appendix B Scaling freezing quintessence
In [30] it was demonstrated that in appropriate variables, resulting in a global 3D state space setting, scaling freezing quintessence is associated with a single attractor orbit that constitutes the unstable manifold of an isolated fixed point. The situation and key attractor mechanism is thereby the same as for tracking quintessence. An open set of nearby orbits to the (hyperbolic) saddle fixed point is pushed toward the attractor orbit by a 2D invariant unstable manifold boundary set, for details, see [30] and [32]. The difference with tracking quintessence is that scaling freezing quintessence arises from potentials that have an asymptotic exponential behaviour (that is sufficiently steep).
@@
$V=V_{-}e^{-\lambda_{-}\phi}+V_{+}e^{-\lambda_{+}\phi},\qquad\lambda_{-}>\lambda_{+},$ (82)
@@
$\Gamma=1+\frac{(\lambda-\lambda_{-})(\lambda_{+}-\lambda)^{2}}{\lambda^{2}}.$ (83)
@@
Note that (83) is regular at $\lambda=\lambda_{-}$ for the potential with two exponential terms and that
$\Gamma_{1}=\left(1-\frac{\lambda_{+}}{\lambda_{-}}\right)^{2},\qquad\Gamma_{2}=\frac{2}{\lambda_{-}}\left(1-\left(\frac{\lambda_{+}}{\lambda_{-}}\right)^{2}\right)$ (86)
in this case
@@
Note that $\lim_{N\rightarrow-\infty}w_{\phi}=0$, which leads to that asymptotically toward the past $\rho_{\phi}\propto\rho_{\mathrm{m}}\propto a^{-3}$, while $w_{\phi}$ decreases toward the future, hence the name scaling freezing.

### v3

@@ A.2 Tracking quintessence
Here we consider the tracking quintessence approximations given by Chiba (2010) [57], which in turn relied on earlier work by Watson and Scherrer (2003) [63], for the case $\Gamma=\mathrm{constant}$, and, in particular, $\Gamma=\mathrm{constant}>1$, which yields the inverse power law potential. This work is based on a linearized equation for a perturbation of $w_{\phi}$ with respect to its asymptotic value $w_{\infty}$. To obtain an improved approximation for the linearized solution for $w_{\phi}$ Chiba [57] considered a series expansion in $e^{-3wN}$, when expressed in $e$-fold time and the present notation. This is the same series expansion quantity as in the present paper, but in contrast to Chiba [57] we used it to produce approximations for the full non-linear equations; thus, the expansions in the present work coincide with those in [63] and [57] for the quantities that have been computed in those two references up to first order but not for higher orders since Chiba [57] only deals with the linearized equation for $w_{\phi}$. Moreover, there is translation freedom in $N$. In the present work this freedom was fixed by the $[1/1]_{\Omega_{\phi}}$ Padé approximant for $\Omega_{\phi}$, which resulted in an expansion in $T=\Omega_{\phi 0}e^{-3w_{\infty}N}/(1-\gamma\Omega_{\phi 0})$; Chiba [57] on the other hand fixed this freedom by, in effect, setting $w_{\phi}=w_{\infty}$, to obtain the following approximation for $\Omega_{\phi}$:
@@
which corresponds to $w$CDM, while the series expansion in $\tilde{T}$ for the linearized equation for $w_{\phi}$ resulted in
@@
Note that this is an expansion that is designed to improve the solution for the linearized equation for $w_{\phi}$, not the actual solution of the non-linear equation for $w_{\phi}$. The improved linear approximation only improves the actual solution to the extent the linear approximation describes that solution, which, fortunately, to some extent is the case for the inverse power law potential, see e.g., Figure 1 in [30]. However, not surprisingly, these approximations yield poorer accuracy than the present ones that are based on approximating the full non-linear equations.
@@ Appendix B Scaling freezing quintessence
In [36] it was demonstrated that in appropriate variables, resulting in a global 3D state space setting, scaling freezing quintessence is associated with a single attractor orbit that constitutes the unstable manifold of an isolated fixed point. The situation and key attractor mechanism is thereby the same as for tracking quintessence. An open set of nearby orbits to the (hyperbolic) saddle fixed point is pushed toward the attractor orbit by a 2D invariant stable manifold boundary set, for details, see [36] and [37]. The difference with tracking quintessence is that scaling freezing quintessence arises from potentials that have an asymptotic exponential behaviour that is sufficiently steep.
@@
$V=V_{-}e^{-\lambda_{-}\phi}+V_{+}e^{-\lambda_{+}\phi},\qquad\lambda_{-}>0,\qquad\lambda_{-}>\lambda_{+},\qquad V_{\pm}>0,$ (100)
@@
$\Gamma=1+\frac{(\lambda_{-}-\lambda)(\lambda-\lambda_{+})}{\lambda^{2}}.$ (101)
@@
Note that (101) is regular since $\lambda>\lambda_{-}>0$ for the potential with two exponential terms and that
$\Gamma_{1}=-\left(\frac{\lambda_{-}-\lambda_{+}}{\lambda_{-}^{2}}\right),\qquad\Gamma_{2}=\frac{2(\lambda_{-}-2\lambda_{+})}{\lambda_{-}^{3}}$ (104)
in this case.
@@
Note that $\lim_{N\rightarrow-\infty}w_{\phi}=0$, which leads to that asymptotically toward the past $\rho_{\phi}\propto\rho_{\mathrm{m}}\propto a^{-3}$, hence the name scaling freezing, while $w_{\phi}$ decreases toward the future.